# Automatic Link Selection in Multi-Channel Multiple Access with Link Failures

### v1

@@ VII Simulations
For simulations, we consider four scenarios with $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+x_{i})$, $n=5,m=3$, and $T=10^{5}$. In each scenario, the values of $q_{i,j}$ change mid-way through the simulation (at $T=5\times 10^{4}$). In Scenario 1 (Figure 1-Left-Top), certain user-channel links are turned off at $T=5\times 10^{4}$. In Scenario 2 (Figure 1-Right-Top), certain user-channel links that were turned off before, turn on at $T=5\times 10^{4}$. In Scenarios 3 (Figure 1-Left-Bottom) and 4 (Figure 1-Right-Bottom), the values of $q_{i,j}$ for certain user-channel links change at $T=5\times 10^{4}$. In each scenario, the algorithms are not told about the change. For each figure, for each algorithm, we plot $\phi\left(\frac{1}{t}\sum_{\tau=1}^{t}\boldsymbol{X}(\tau)\right)$ for $t\leq T/2$ and $\phi\left(\frac{1}{t-T/2}\sum_{\tau=T/2+1}^{t}\boldsymbol{X}(\tau)\right)$ for $t>T/2$ vs $t$. We also plot the optimal objective value of (P2), for reference. Note that in the first half of the simulation, both algorithms converge to the optimal solution in all scenarios, although the convergence of Algorithm 3 is faster. However, to show the importance of Algorithm 1, we observe the second half of the simulation. Although Algorithm 3 achieves superior performance in the first half of the simulation, it does not perform well in the second half except in Scenario 2. In contrast Algorithm 1 achieves similar performance in both halves in all scenarios. This is due to the adaptiveness of Algorithm 1.
Figure 1: Top-Left: Scenario 1. Top-Right: Scenario 2. Bottom-Left: Scenario 3. Bottom-Right: Scenario 4.
@@ VIII Conclusions
This paper focused on the problem of designing algorithms for automatic link selection in multi-channel multiple access with link failures. In particular, we solved a network utility maximization problem with matching constraints and bandit feedback on link failures. We considered two algorithms, where the first algorithm has slower convergence and is adaptive, and the second algorithm has faster convergence and is not adaptive.
Appendix A Proof of Lemma 2
@@
3) First, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}=0$. Then we have $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=0$. Also, due to the scaling, we have $s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}\geq 0$. Hence, we have $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=0\leq s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$ and we are done. Next, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}>0$. Notice that
$\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$ $=\frac{\lVert(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
$=\frac{\lVert\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
@@
where (a) follows since $\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\geq 0$, and $\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1}\geq 0$, as a result of which $(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\geq 0$.
4) Denote by $\mathcal{U}=\left\{i\in[s]:\sum_{j=1}^{s}P_{i,j}>1\right\}$ and $\mathcal{U}^{c}=[s]\setminus\mathcal{U}$. Notice that
$\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}+\lVert\boldsymbol{P}\rVert_{1}-s$
$=\sum_{i=1}^{s}\left\lvert\sum_{j=1}^{s}P_{i,j}-1\right\rvert+\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}-s$
@@
$\geq 2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\left(\sum_{j=1}^{s}\left\lvert\sum_{i=1}^{s}P_{i,j}-1\right\rvert\right)$
@@
$\lVert\boldsymbol{P}-\boldsymbol{Q}\rVert_{1}$ $\leq\lVert\boldsymbol{P}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
$=_{(a)}\lVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
@@
where (a) follows since $\boldsymbol{P}^{{}^{\prime\prime}}\leq\boldsymbol{P}$ due to the scaling, (b) follows due to part 3, and (c) follows from (54).
Appendix B Proof of Lemma 20
Notice that from the queuing equation (41), we have
$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)]$
for all $i\in[n]$. Summing the above for $i\in[n]$, we have
where the last equality follows since $\boldsymbol{Y}(t),\boldsymbol{\gamma}(t),\boldsymbol{Q}(t)$ are $\mathcal{H}(t)$-measurable. Taking expectations, we have the result.

### v2

@@ VII Simulations
We consider eight simulation scenarios. All experiments were conducted on a virtual machine equipped with an Intel Xeon 2.20 GHz CPU (2 virtual CPUs). The algorithms were implemented in Python. In the first four scenarios, we use the smooth utility function $\phi(\mathbf{x})=\sum_{i=1}^{n}\log(1+b_{i}x_{i})$, where $b_{i}>0$ are fixed constants. In scenarios 5 and 6, we consider the nonsmooth utility function $\phi(\mathbf{x})=a\sum_{i=1}^{n}x_{i}+b\min(x_{1},x_{2},\dots,x_{n})$, where $a,b>0$, are fixed constants. In the last two scenarios, we use the utility function $b\min(x_{1},x_{2},\dots,x_{n})$, where $a,b>0$, is a fixed constant.
We run each algorithm for $T=10^{6}$ time slots. The channel success probabilities are changed once at time $T_{0}=5\times 10^{5}$, in order to evaluate the adaptiveness of the algorithms. The algorithms are neither restarted nor informed of this change. To examine performance before and after the change point, we plot the evolution of the objective applied to empirical averages of the observed vectors $\boldsymbol{X}(t)$.
We compute
$\Phi(t)=\begin{cases}\phi\!\left(\frac{1}{t}\sum_{\tau=1}^{t}\mathbf{X}(\tau)\right),&\text{ if }t<T_{0}\\ \phi\!\left(\frac{1}{t-T_{0}+1}\sum_{\tau=T_{0}}^{t}\mathbf{X}(\tau)\right)&\text{ if }t\geq T_{0}\end{cases}$ (98)
We then plot $\Phi(t)$ as a function of $t$. This representation allows us to separately examine the convergence behavior before the change and the transient adaptation after the change. In each plot, we also include the optimal utility value obtained by solving Problem (P1) using the corresponding channel success probabilities. Figure 1 denotes the plots. In each plot, the y-axis is zoomed in to clearly highlight the performance differences among the algorithms. If the curve corresponding to a particular algorithm is not visible in the second half of the plot, it indicates that the algorithm has not adapted sufficiently to reach the new optimal value.
(a) Scenario 1
(b) Scenario 2
(c) Scenario 3
(d) Scenario 4
(e) Scenario 5
(f) Scenario 6
(g) Scenario 7
(h) Scenario 8
Figure 1: Evolution of the objective $\Phi(t)$ defined in (98). A vertical dashed line marks $T_{0}=5\times 10^{5}$, the point at which the success probabilities change.
In Scenarios 1 and 2, we consider a multi-channel setting with $n=4$ and $m=3$. In Scenario 1, the success probabilities $q_{i,j}$ in the two halves of the simulation are chosen randomly. In Scenario 2, we interchange these $q_{i,j}$ values. Here, we compare the Adaptive MAC (Algorithm 1), Adaptive MAC.CF (Algorithm 4), and UCB-MAC (Algorithm 4) algorithms. Notice that the adaptive algorithms have decent performance in the first half of the simulation. Although UCB-MAC has faster convergence in the first half, it does not adapt to changes in success probabilities in the second half. An interesting observation is that, despite the stronger worst-case guarantee of the Adaptive MAC algorithm, the Adaptive MAC.CF algorithm appears to outperform it in simulation.
In Scenarios 3 and 4, we consider a single-channel setting with $n=4$. We assume that some links may become unavailable without the controller’s knowledge; in that case, the corresponding $q_{i,j}$ is set to zero. When a link is available, its success probability is $q_{i,j}=0.4$. Within each half of the simulation horizon, the availability status of each link remains fixed. Here, for the adaptive algorithm, we use the Single-Channel Adaptive MAC (Algorithm 3). We also compare with the Algorithms introduced in [2] (Algorithm 1 of [2], and Frank-Wolfe-based Algorithm 3 of [2]). In this case, the adaptation capability of the adaptive MAC algorithm can be clearly seen. In fact, in Scenario 3, the UCB-MAC algorithm performs extremely poorly in the second half of the simulation, resulting in an objective value that drops close to zero.
For Scenario 5 and Scenario 6, we again consider a single channel scenario. Recall that the utility function in this case is $\phi(\boldsymbol{x})=a\sum_{i=1}^{n}x_{i}+b\min(x_{1},x_{2},\dots,x_{n})$. We compare Single-Channel Adaptive MAC, UCB-MAC, and Algorithm 1 of [2]. Notice that even for the nonsmooth utility, the adaptive algorithm adapts fast in the second half of the simulation.
In Scenarios 7 and 8, we consider the single-channel setting with utility $\phi(\boldsymbol{x})=b\min(x_{1},x_{2},\dots,x_{n})$. We also plot the simple renewal mechanism described in Section VI-C. In these scenarios, UCB-MAC also adapts in the second half of the simulation. Interestingly, we observe that UCB-MAC does not rely on explicit UCB-based estimation in this case. Instead, it effectively implements a mechanism that closely resembles the renewal mechanism.
In Table II, we report the average runtime per time slot of the algorithms under different scenarios. We observe that Adaptive MAC.CF has faster iterations than Adaptive MAC. However, the UCB-based non-adaptive algorithm outperforms both adaptive algorithms in terms of runtime. In the single-channel scenario, the adaptive algorithm achieves runtimes comparable to those of the UCB MAC algorithm. Finally, the UCB-based non-adaptive algorithm outperforms the algorithms proposed in [2] in terms of empirical runtime complexity.
TABLE II: Average runtime per time slot(in milliseconds) of each algorithm under each scenario. For the Adaptive MAC algorithm, we also report (in parenthesis) the average number of Sinkhorn iterations per main iteration.
Algorithm Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 Scenario 6 Scenario 7 Scenario 8
Adaptive 3.878 3.849 - - - - - -
MAC (12.982) (12.054) - - - - - -
Adaptive 2.655 2.433 - - - - - -
MAC.CF
UCB MAC 0.207 0.211 0.157 0.171 0.0682 0.07 0.0651 0.0655
Single
Channel - - 0.242 0.27 0.0729 0.0744 0.0706 0.0715
MAC
BwCR
Algorithm 1 - - 9.224 9.556 0.0691 0.0773 0.0672 0.0655
[2]
BwCR
Algorithm 3 - - 0.17 0.171 - - - -
[2]
@@ VIII Conclusions
This paper focuses on the problem of designing algorithms for automatic link selection in multi-channel multiple access systems with transmission failures. In particular, we solve a network utility maximization problem with matching constraints and bandit feedback on transmission outcomes. We propose two adaptive algorithms: the first achieves faster convergence, while the second admits a more efficient implementation. We also study several special cases with additional structural properties. Extending the framework to adversarial settings and incorporating more general combinatorial constraints are directions for future work.
Appendix A SINKHORN function
We introduce SINKHORN proceure on positive $s\times s$ matrices.
[…]
Initialize $\boldsymbol{R}^{1}=\boldsymbol{R}$.
For each iteration $\tau\in\{1,2\dots,\}$
  1. 1.
    Row normalization: If $\tau$ is even set $\boldsymbol{R}^{\tau+1}$ as
    $R_{i,j}^{\tau+1}=\frac{R_{i,j}^{\tau}}{\sum_{k=1}^{s}R_{i,k}^{\tau}}$ (99)
  2. 2.
    Column normalization: Else set $\boldsymbol{R}(\tau+1)$ as
    $R_{i,j}^{\tau+1}=\frac{R_{i,j}^{\tau}}{\sum_{k=1}^{s}R_{k,j}^{\tau}}$ (100)
Given $\boldsymbol{X}\in\mathbb{R}^{s\times s}$, and $\boldsymbol{Q}\in(0,\infty)^{s\times s}$, SINKHORN studied in [24] is an iterative process to solve problems of the type
$\min_{\boldsymbol{P}\in\mathcal{S}^{\text{doub}}}\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}P_{i,j}+D(\boldsymbol{P}\lVert\boldsymbol{Q})$ (101)
We provide the following lemma adapted from [24]
Lemma 24
Given $\epsilon>0$, running SINKHORN($\boldsymbol{R}$), where $\boldsymbol{R}$ is defined by $R_{i,j}=Q_{i,j}\exp(-X_{i,j})$ for sufficiently many iterations (say $U$ iterations) generates a matrix $\boldsymbol{R}^{U}$ such that
$\exp(-\epsilon)\leq\frac{R^{U}_{i,j}}{P^{*}_{i,j}}\leq\exp(\epsilon)$ (102)
where $\boldsymbol{P}^{*}$ is the solution of Problem (101).
Appendix B Finding $\boldsymbol{\hat{P}}(t+1)$ in Problem (26)
Finding $\boldsymbol{\hat{P}}(t+1)$ can be summarized under the following steps.
  1. 1.
    Define matrix $\boldsymbol{R}\in(0,\infty)^{s\times s}$ by $R_{i,j}=P_{i,j}(t)\exp(\eta Q_{i}(t)\hat{S}_{i,j}(t))$
  2. 2.
    Run SINKHORN($\boldsymbol{R}$) for $U$ iteration to get $\boldsymbol{R}^{U}$ such that
    $e^{-\theta/(3s)}\leq\frac{R^{U}_{i,j}}{\tilde{P}_{i,j}(t+1)}\leq e^{\theta/(3s)}$ (103)
    This is possible due to Lemma 24. See Appendix A for details.
  3. 3.
    Set $\boldsymbol{\hat{P}}(t+1)=\text{ROUND}(\boldsymbol{R}^{U})$. See Section V-A for details.
We have the following lemma that establishes the required property of $\boldsymbol{\hat{P}}(t+1)$.
Lemma 25
We have that $\boldsymbol{\hat{P}}(t+1)$ found above is a $(\theta/s)$-approximation of $\boldsymbol{\tilde{P}}(t+1)$
Proof:
First, notice that for each $i\in[s]$
$\sum_{k=1}^{s}R_{i,k}^{U}\leq\sum_{k=1}^{s}\tilde{P}_{i,j}(t+1)e^{\theta/(3s)}=e^{\theta/(3s)},$ (104)
where the inequality follows from (103), and the equality follows since $\boldsymbol{\tilde{P}}(t+1)$ is doubly stochastic by definition. Similarly, for each $j\in[s]$, we have $\sum_{k=1}^{s}R_{k,j}^{U}\leq e^{\theta/(3s)}$.
First, from Lemma 12, we have that $\boldsymbol{\hat{P}}(t+1)\in\mathcal{S}^{\text{doub}}$. Also, from the description of the ROUND function in Section V-A, notice that
$\hat{P}_{i,j}(t+1)$ $\geq\frac{R_{i,j}^{U}}{(\sum_{k=1}^{s}R_{i,k}^{U})(\sum_{k=1}^{s}R_{k,j}^{U})}\geq_{(a)}\frac{R_{i,j}^{U}}{e^{2\theta/(3s)}}\geq\frac{\tilde{P}_{i,j}(t+1)}{e^{\theta/s}}$ (105)
where (a) follows from the previous two inequalities, and (b) follows from (103). Hence, we are done. ∎
Appendix C Proof of Lemma 12
@@
3) We first prove the following claim
Claim: We have that $\lvert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}\leq s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
Proof:
First, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}=0$. Then we have $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=0$. Also, due to the scaling, we have $s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}\geq 0$. Hence, we have $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=0\leq s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$ and we are done. Next, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}>0$. Notice that
$\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$ $=\frac{\lVert(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}=\frac{\lVert\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
@@
where (a) follows since $\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\geq 0$, and $\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1}\geq 0$, as a result of which $(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\geq 0$. ∎
Now we prove part 3. Denote by $\mathcal{U}=\left\{i\in[s]:\sum_{j=1}^{s}P_{i,j}>1\right\}$ and $\mathcal{U}^{c}=[s]\setminus\mathcal{U}$. Notice that
$\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}+\lVert\boldsymbol{P}\rVert_{1}-s=\sum_{i=1}^{s}\left\lvert\sum_{j=1}^{s}P_{i,j}-1\right\rvert+\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}-s$
@@
[…]
@@
$\lVert\boldsymbol{P}-\boldsymbol{Q}\rVert_{1}$ $\leq\lVert\boldsymbol{P}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=_{(a)}\lVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
@@
where (a) follows since $\boldsymbol{P}^{{}^{\prime\prime}}\leq\boldsymbol{P}$ due to the scaling, (b) follows due to the claim, and (c) follows from (107).
Appendix D Finding $\boldsymbol{\tilde{P}}(t+1)$ in Problem (55)
We only consider the case where $t$ is even. The case where $t$ is odd can be solved similarly. Notice that we can separately solve for each column of $\boldsymbol{\tilde{P}}(t+1)$. To solve for the $j$-th column of $\boldsymbol{\tilde{P}}(t+1)$, we define $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{s}$, where
$\text{s.t. }\boldsymbol{p}\in\Delta^{s}_{\varepsilon},$
It should be noted that (P3) has a classic structure that is solved in [27]. In particular, we define $\boldsymbol{z}$, where $z_{i}=y_{i}\exp(x_{i})$. First, assume that $\boldsymbol{z}$ is sorted in the increasing order. Then it can be shown that there exists $i\in[0:s-1]$ such that the vector $\boldsymbol{u}^{i}\in\mathbb{R}^{s}$ given by,
[…]